OpenStudy (hihi67):

2*a^4*4*a^2= 3*b^5*6*b= 4*b^4*3/b^3= 15*x^6*3/x^2= I solved these, but want to know if i have the right answer.

OpenStudy (hihi67):

for the first one i got 8*a^6

OpenStudy (hihi67):

#2 18*b^5

OpenStudy (hihi67):

12*b^1

OpenStudy (hihi67):

or 12*b

OpenStudy (hihi67):

and last 45*x^4

OpenStudy (hihi67):

what about you?

OpenStudy (cathyangs):

I got 8a^6 12b and 5x^4

OpenStudy (hihi67):

there are 4 problems though

OpenStudy (cathyangs):

oops, Missed the second one. I got 18b^6

OpenStudy (anonymous):

I got 8a^6 12b and 45x^4.. Im sure all are correct..

OpenStudy (alfie):

First one is fine. Second one it's 18b^6 The other ones, you have to write them through the equation symbol down here otherwise we can't understand properly the equation (use parenthesis! :) )

OpenStudy (hihi67):

so on the second one is it b^6 because b=b^1?

OpenStudy (anonymous):

sry,... me too missed the second one... ans is 18b^6

OpenStudy (cathyangs):

yep, and 5+1=6

OpenStudy (hihi67):

yep

OpenStudy (hihi67):

thanks!

OpenStudy (cathyangs):

ahahaXD the irony (ish). we both missed the second one

OpenStudy (anonymous):

\[\left\{8 a^6,18 b^6,12 b,45 x^4\right\} \]